A study found that the mean systolic blood pressure was 106 mm Hg. Assume that
systolic blood pressure follows a normal distribution with a standard deviation
of 8.
A.) What is the probability that a randomly selected systolic blood pressure
is between 98 and 122 mm Hg?
B.) Find the 97.5th percentile of systolic blood pressures. Interpret the
97.5th percentile.

Answer by ewatrrr(24785)

 
Hi,
μ = 106
σ = 8
A) z = -8/8 = 1 and z = 16/8 = 2
P(-1 > z < 2) = .9883 - .1587 = .8186 or 81.86%
NORMSDIST(2) NORMSDIST(-1)
B) 97.5 percentile P = .975 the z = 1.96
1.96 *8 + 106 = = 121.68 or rounded up 122mm Hg
97.5 % of the people have a systolic blood pressure less than 122